# GINTRIP: Interpretable Temporal Graph Regression using Information bottleneck and Prototype-based method

### v1

@@ III Experiments
We evaluate the performance of GINTRIP in terms of forecasting metrics and explainability on three distinct real-world traffic datasets. In terms of explainability metrics, the fidelity and sparsity are defined as[16]:
$\small\text{Fidelity}_{+/-}(\%)=\frac{100}{\lfloor T/W\rfloor}\sum_{i=1}^{\lfloor T/W\rfloor}\left(|f(G_{i})-f(G_{i}^{+/-})|\right)$
$f$ represents the trained predictive spatio-temporal function. $G_{i}^{-}$ and $G_{i}^{+}$ refer to the masked spatial-temporal complementary subgraph and the selected subgraph structure for $i^{th}$ $G_{in}$, respectively. It directly measures the quality of the learned prototypes on the targeted task. $k$ represent the number of important nodes in the original graph, denoted as Sparsity. Node importance is determined based on the assignment probability explained in Section II-D.
@@ III-A Quantification analysis
We conducted our experiments on real-world traffic datasets. Specifically, PeMS04, PeMS07, and PeMS08 [22, 23] were chosen as graph-based traffic datasets.
@@
In Table I, the traffic forecasting metrics are presented. The results indicate that self-explainablity promotes the predictivity capacity. In contrast to [14], our extracted subgraphs are time-aware, which eases the interpretation process. Indeed, it dispenses with the post-processing of temporal and spatial subgraphs proposed in [14].
TABLE I: Traffic predictions benchmark, B.B. is the abbreviation for block box models which are not interpretable
@@
To verify our interpretability, we compare the fidelity and sparsity with STExplainer[14], which is presented in Fig.2. Our fidelity score($F_{+}$) is superior to the previous state-of-the-arts.
Figure 2: Fidelity over Sparsity. High fidelity is an indicator of interpretability
@@ III-B Visualizations
Fig. 3 illustrates the interpretability of the learned model by presenting extracted subgraphs at various sparsity thresholds ($k$). These subgraphs, derived from the Pems04 dataset, exhibit consistent spatio-temporal patterns across the selected subgraph nodes regarding different $k$ values. This consistency underscores GINTRIP’s ability to extract robust, time-aware subgraphs using the IB technique and learned prototypes.
@@
Figure 3: Positive over the negative fidelity based on the two selected subgraphs regarding the sparsity (k)
IV Conclusion
@@
In this study, we emphasize the significance of explainability in spatio-temporal graph neural networks. To address this, we propose a novel framework called GINTRIP that not only predicts future spatio-temporal signals accurately but also provides interpretable time-aware subgraphs. Our framework incorporates the prototype learning approach which employs variational approximation for tractability. Additionally, we introduce a unified Temporal-GNN that provides intrinsically explainable and robust spatio-temporal representations. Through comprehensive experiments, we examine the hypothesis that self-explainability causes predictability and robustness. Our results surpass existing state-of-the-art baselines in both predictive accuracy and explainability. In future research, we plan to investigate interpretability in hypergraphs.
@@ Acknowledgment
This work is funded by Innoviris within the research projects TORRES. AR and LDC acknowledge support from the Belgian Science Policy Office (BELSPO) under contract number B2/233/P2/PRECIP-PREDICT and through the FED-tWIN programme (Prf-2020-017).

### v2

@@ III Experiments
We evaluate the performance of GINTRIP in terms of forecasting metrics and explainability on three distinct real-world traffic and crime datasets. In terms of explainability metrics, the fidelity is defined as [21]:
$\small\text{Fidelity}^{+/-}=\frac{1}{\lfloor T/W\rfloor}\sum_{i=1}^{\lfloor T/W\rfloor}\left(|f(G_{i})-f(G_{i}^{+/-})|\right)$ (17)
$f$ represents the trained predictive spatio-temporal function. $G_{i}^{-}$ and $G_{i}^{+}$ refer to the selected subgraph structure and the complementary structure of the selected subgraph in G respectively, for $i^{th}$ $G_{in}$. It directly measures the quality of the learned prototypes on the targeted task. Also, in Figure 3, density, $k$, denotes the ratio of important nodes (the number of nodes in subgraph structures) to total nodes in the original graph. Node importance is determined based on the assignment probability explained in Section II-D.
@@ III-A Quantification analysis
We conducted our experiments on real-world traffic datasets. Specifically, PeMS04, PeMS07, and PeMS08 [27] were chosen as graph-based traffic datasets.
@@
The traffic forecasting metrics are presented in Table I. The results indicate that self-explainablity promotes the predictivity capacity. In contrast to [19], our extracted subgraphs are time-aware, which eases the interpretation process. Indeed, it dispenses with the post-processing of temporal and spatial subgraphs proposed in [19].
TABLE I: Traffic predictions benchmark, $B.$ is the abbreviation for block box models which are not interpretable
@@
To validate GINTRIP’s generalizability, we conducted comprehensive experiments across multiple crime datasets, NYC and Chicago[30]. The results are presented in TableII. Consistently, GINTRIP outperforms specialized crime prediction methods like DMSTGCN [31] attributed to its interpretability. Based on the superior or same performance, GINTRIP holds up the statement that self-explainability causes predictability and robustness.
TABLE II: Crime prediction results on NYC and Chicago datasets
Model NYC Crime CHI Crime
MAE MAPE MAE MAPE
ST-SHN [30] 0.9280 0.5373 1.0689 0.5116
DMSTGCN [31] 0.9293 0.5485 1.0736 0.5175
STExplainer [19] 0.9095 0.5154 1.0701 0.5143
GINTRIP (ours) 0.8912 0.5016 1.055 0.5117
To verify our interpretability, we compare the fidelity and $K$ with STExplainer[19], which is presented in figure 2. Our $Fidelity^{+}$ score is superior to the previous state-of-the-art. In the figure 2, the x-axis is the density parameter. So, mostly, with the different degrees of the subgraph density, the $Fidelity^{+}$ of our model is better than STExplainer[19], implying superior interpretability.
Figure 2: Positive fidelity over k. High fidelity is an indicator of interpretability.
@@ III-B Visualizations
figure 3 illustrates the interpretability of the learned model by presenting extracted subgraphs at various density thresholds ($k$). These subgraphs, from the Pems04 dataset, exhibit consistent spatio-temporal patterns across the selected subgraph nodes regarding different values of $k$. This consistency underscores GINTRIP’s ability to extract robust, time-aware subgraphs using the IB technique and learned prototypes. Our model detects soundly enriched subgraphs such that, by increasing their density, e.g. $k$, the inverse relationship between these two metrics can be observed.
@@
Figure 3: Positive over the negative fidelity based on the two selected subgraphs regarding the k.
As a significant achievement, as opposed to STExplainer[19], which finds disjointed spatial and temporal subgraphs causes hampering the interpretation of the temporal interpretability, we introduce a unified subgraph that jointly encoded spatial and temporal interpretability. To bolster this concept, we plotted learned subgraphs by network unfolded in time in figure 4 for Pems04 dataset. For the first three consecutive prediction windows, our model did not select the edge $47\rightarrow 36$, while in the input graph, they are adjacent. The underlying reason would be the temporal cross-correlation of the corresponding time series is not expressive enough. The selection probabilities determined by the model truly are proportional to the pairwise correlation.
Figure 4: Explanation of the learned prototypes for a spatio-temporal prediction example. The y-axis denotes traffic speed.
(a) Chicago crime rate prediction benchmark.
(b) New York City crime rate prediction benchmark.
Figure 5: Crime rate prediction benchmarks for Chicago and New York City.
III-C Implementation Details
[…]
IV Ablation Study
IV-1 Dense Area Analysis
The ablation study has been conducted for the prediction in the dense area for both the NYC and Chicago datasets, which are depicted in figure 5. Consistently, in terms of prediction metrics, our model outperforms STExplainer[30].
IV-2 Hyperparameter Analysis
Figure 6: Hyperparameter importance ablation study
[…]
IV-3 Prototype Module Importance
PeMS04 PeMS07
Run $downarrow MAE\downarrow RMSE\downarrow$
W/O Prototype 18.85 29.26 20.28 33.71
Ours 18.62 29.02 20.12 33.53
TABLE III: Ablation study for the prototype module importance
Table III presents the ablation study conducted to assess the contribution of the proposed prototype module. The first row reports the performance when the module is disabled, resulting in higher errors across both datasets. In contrast, incorporating the module consistently improves MAE and RMSE on PeMS04 and PeMS07, demonstrating its effectiveness. These results highlight the necessity of the prototype module in improving predictive accuracy and validate its role as a key component of our framework.
IV-4 Pseudo-label Robustness
Method Parameter MAE
Quantile 0.80 18.66
K-Means [34] 4 18.912
DB-SCAN [35] - 18.928
K-Means [34] 3 18.915
Quantile 0.70 18.75
Quantile (ours) 0.90 18.62
TABLE IV: Sensitivity analysis of GINTRIP regarding quantile threshold on Pems04 dataset.
Figure 7: K-Means metric ablation study on Pems04 dataset.
As described in Section II-A, we use a quantile-based function to determine pseudo-labels. In Table IV, we ablate different threshold values. In addition, inspired by [35], we implement K-means and DBSCAN to estimate the threshold in an unsupervised manner; the resulting clusterings are shown in Figure 7. Overall, the different thresholding strategies yield comparable performance.
V Conclusion
@@
In this study, we address the need for explainability in spatio-temporal graph neural networks by introducing a novel framework called GINTRIP. This method not only predicts future spatio-temporal signals accurately but also provides interpretable time-aware subgraphs. Our framework incorporates the prototype learning approach which employs variational approximation for tractability. Additionally, we introduce a unified Temporal-GNN that provides intrinsically explainable and robust spatio-temporal representations. Through comprehensive experiments, we examine the hypothesis that self-explainability causes predictability and robustness. Our results surpass existing state-of-the-art baselines in both predictive accuracy and explainability. In future research, we plan to investigate interpretability in hypergraphs.
V-A Limitations and Future Work
While this paper provides insights into spatio-temporal neural networks, it implicitly assumes a static connectivity structure, overlooking the dynamics of changing connections. Future research could enhance our framework by incorporating dynamic causal models, allowing the model to adapt to variations in causal relationships over different connection times.
Appendix A Summary of Notations
For clarity, Table V summarizes the key notations used throughout this paper.
TABLE V: Summary of Notations
Symbol Definition
Graph Structure
$G_{in}$ Input temporal graph
$V_{G_{in}}$ Set of nodes in the input graph, $V_{G_{in}}=\{V_{1},V_{2},\ldots,V_{N}\}$
$E_{G_{in}}$ Set of edges in the input graph
$A_{G_{in}}\in\mathbb{R}^{N\times N}$ Adjacency matrix of the input graph
NN Total number of nodes in the graph
$G_{sub}$ Extracted subgraph from $G_{in}$ via the IB principle
$G_{p}$ Set of learnable prototypes, $G_{p}=\{\mathbf{z}_{G_{p}}^{1},\mathbf{z}_{G_{p}}^{2},\ldots,\mathbf{z}_{G_{p}}^{M}\}$
Temporal and Feature Variables
$X_{t}\in\mathbb{R}^{N\times D}$ Dynamic node feature matrix at timestamp $t$
$D$ Dimensionality of node features
$T$ Length of the observed time series history
$T^{\prime}$ Length of the forecast horizon
$W$ Sliding window length ($W\leq T$)
$\bm{x}\in\mathbb{R}^{W}$ Model input: sliding window $y_{T-W+1},\ldots,y_{T}]$
$\bm{Y}$ Prediction series (output)
$Y_{reg}$ Ground-truth regression target
$\hat{Y}_{reg}$ Predicted regression output
$Y_{cls}$ Classification target (pseudo-labels)
Latent Representations
$h_{i}$ Latent representation of node $i$ from the GNN encoder
$z_{i}$ Final representation of node $i$ in the extracted subgraph
$Z_{sub}$ Pooled embedding of the subgraph $G_{sub}$
$\mathbf{z}_{G_{p}}^{m}$ The $m$-th learnable prototype vector
$Z_{G_{p}}$ Concatenation of all prototype vectors $\{\mathbf{z}_{G_{p}}^{1},\ldots,\mathbf{z}_{G_{p}}^{M}\}$
Prototype and Classification Variables
$M$ Total number of prototypes ($M=K\times J$)
$K$ Number of pseudo-classes (e.g., congested/non-congested)
$J$ Number of prototypes allocated per class
$T_{i}$ Threshold for classifying node $V_{i}$, defined as $Q_{0.1}(X_{:,i,:})$
$S^{m}$ Similarity score between $Z_{sub}$ and the $m$-th prototype $\mathbf{z}_{G_{p}}^{m}$
Subgraph Extraction Variables
$p_{i}$ Probability of node $i$ being included in the subgraph $G_{sub}$
$\lambda_{i}^{node}$ Bernoulli sample $\lambda_{i}^{node}\sim\text{Bernoulli}(p_{i})$ for node selection
$\epsilon$ Gaussian noise, $\epsilon\sim\mathcal{N}(\mu_{h_{i}},\sigma_{h_{i}}^{2})$
$\mu_{h_{i}}$ Mean of the latent representation $h_{i}$
$\sigma_{h_{i}}$ Standard deviation of the latent representation $h_{i}$
$P\in\mathbb{R}^{|V_{G}|\times 2}$ Node probability assignment matrix for connectivity loss
$k$ Subgraph density: ratio of selected nodes to total nodes
Information-Theoretic Terms
$I(X;Y)$ Mutual information between random variables $X$ and $Y$
$H(X)$ Entropy of random variable $X$
$\beta$ Lagrange multiplier controlling compression-prediction trade-off in IB
Functions and Operators
$\gamma(\cdot,\cdot)$ Similarity function (e.g., dot product) between $Z_{sub}$ and $G_{p}$
$q_{\phi}(\cdot)$ Variational approximation to posterior $p(Y_{cls}|\gamma(G_{sub},G_{p}))$
$q_{\theta}(\cdot)$ Single-layer perceptron estimating $Z_{G_{p}}$ from $Z_{sub}$
$f(\cdot)$ Trained spatio-temporal predictive function
$\|\cdot\|_{F}$ Frobenius norm
$\widehat{(\cdot)}$ Row normalization operator
Loss Functions and Weights
$\mathcal{L}$ Total loss function
$\mathcal{L}_{reg}$ Regression loss (MSE between $\hat{Y}_{reg}$ and $Y_{reg}$)
$\mathcal{L}_{cls}$ Classification loss (cross-entropy for $Y_{cls}$)
$\mathcal{L}_{sub}$ Upper bound on $I(G_{in};G_{sub})$ for subgraph compression
$\mathcal{L}_{var}$ Variational loss: $\text{MSE}(q_{\theta}(Z_{sub}),Z_{G_{p}})$ for prototype learning
$\mathcal{L}_{con}$ Connectivity loss ensuring subgraph locality (Eq. 9)
$\lambda_{1},\ldots,\lambda_{5}$ Weighting coefficients for each loss component
$\lambda,\lambda^{\prime}$ Constant coefficients in MI lower bounds (Eqs. 12, 13)
Evaluation Metrics
$\text{Fidelity}^{-}$ Prediction change when using only the selected subgraph $G_{i}^{-}$
$\text{Fidelity}^{+}$ Prediction change when using the complement of subgraph $G_{i}^{+}$
@@ Acknowledgment
This work is funded by Innoviris within the research project TORRES under grant number 2022-RDIR-59c. AR and LDC acknowledge support from the Belgian Science Policy Office (BELSPO) under contract number B2/233/P2/PRECIP-PREDICT and through the FED-tWIN program (Prf-2020-017).